OpenStudy (anonymous):

Given x/x^2-9 - x/x^2+6x+9, perform the indicated operation and simplify the result.

OpenStudy (anonymous):

\[\frac{x}{(x-3)(x+3)}-\frac{x}{(x+3)^2}\] \[\frac{x(x+3)-x(x-3)}{(x+3)^2(x-3)}\] \[\frac{x^2+3x-x^2+3x}{(x+3)^2(x-3)}=\frac{6x}{(x+3)^2(x-3)}\]

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

you're welcome